# M4-SAR: A Multi-Resolution, Multi-Polarization, Multi-Scene, Multi-Source Dataset and Benchmark for optical-SAR Fusion Object Detection

### v2

@@ IV-B3 Cross-modal Mamba Interaction Module
R-FCOS [43] $\checkmark$ 31.9M 17.6ms 32.8 79.8 29.7 67.0 78.9 54.8 57.2 26.4 30.2
R-ATSS [65] $\checkmark$ 36.0M 19.1ms 33.2 80.0 29.3 66.6 78.3 54.8 57.0 27.1 30.0
YOLOv5 [16] $\checkmark$ 8.10M 7.20ms 61.4 88.0 56.6 76.4 91.3 91.8 77.6 59.8 52.9
Roi-Trans [33] $\checkmark$ 273M 49.3ms 17.2 80.7 9.1 55.2 80.5 25.6 44.2 27.6 27.1
O-RepPoint [22] $\checkmark$ 36.6M 16.1ms 63.3 90.2 55.8 81.7 88.2 90.1 77.1 58.3 53.9
RTMDet [32] $\checkmark$ 8.86M 11.1ms 34.0 79.8 30.3 61.1 79.1 55.9 56.7 26.5 29.2
YOLOv8 [16] $\checkmark$ 10.0M 6.70ms 63.6 88.6 57.9 80.0 91.3 92.5 79.0 61.6 54.4
PSC [58] $\checkmark$ 31.9M 18.3ms 30.7 79.7 25.9 57.4 78.1 49.4 53.5 26.2 28.5
KFIoU [57] $\checkmark$ 36.2M 66.2ms 16.9 79.4 9.1 50.9 80.4 25.2 43.6 19.4 22.4
YOLOv6 [17] $\checkmark$ 16.0M 6.10ms 57.6 82.1 52.7 67.6 89.4 90.4 73.3 64.9 48.8
YOLOv9 [50] $\checkmark$ 6.40M 14.4ms 61.9 87.7 55.3 74.4 90.0 92.4 76.9 58.9 52.5
YOLOv10 [46] $\checkmark$ 7.50M 7.80ms 65.2 89.9 58.8 82.3 90.8 93.5 80.1 62.8 55.7
YOLOv11 [16] $\checkmark$ 9.74M 6.50ms 64.3 90.0 57.8 80.7 90.9 91.5 79.2 60.5 54.3
LSKNet [24] $\checkmark$ 21.8M 20.4ms 63.3 90.2 55.8 81.7 88.2 90.1 77.1 58.3 53.9
YOLOv12 [42] $\checkmark$ 9.36M 13.4ms 62.7 88.6 57.1 80.1 91.0 92.7 78.7 61.1 54.0
R-FCOS [43] $\checkmark$ 31.9M 17.6ms 23.9 80.1 27.9 60.1 75.8 43.5 51.9 26.4 28.3
R-ATSS [65] $\checkmark$ 36.0M 19.1ms 25.3 80.3 26.8 62.3 77.0 43.9 52.6 24.9 27.8
YOLOv5 [16] $\checkmark$ 8.10M 7.20ms 40.5 85.3 41.0 64.3 85.5 89.9 67.7 44.0 41.3
Roi-Trans [33] $\checkmark$ 273M 49.3ms 9.1 80.8 9.1 35.6 79.0 18.0 38.6 25.2 24.3
O-RepPoint [22] $\checkmark$ 36.6M 16.1ms 26.2 78.6 21.9 53.7 59.9 48.8 48.2 13.9 20.4
RTMDet [32] $\checkmark$ 8.86M 11.1ms 16.2 71.0 21.1 39.3 68.9 39.6 42.7 20.9 22.3
YOLOv8 [16] $\checkmark$ 10.0M 6.70ms 44.9 85.8 42.5 67.9 87.5 89.2 69.6 46.1 43.2
PSC [58] $\checkmark$ 31.9M 18.3ms 22.3 80.0 20.6 51.9 75.7 39.4 48.3 24.5 26.5
KFIoU [57] $\checkmark$ 36.2M 66.2ms 9.1 72.1 9.1 35.3 78.0 22.7 37.7 17.4 19.8
YOLOv6 [17] $\checkmark$ 16.0M 6.10ms 36.5 79.4 32.8 53.5 85.7 86.7 62.4 37.6 36.5
YOLOv9 [50] $\checkmark$ 6.40M 14.4ms 43.9 84.8 38.4 61.5 87.9 88.8 67.6 44.2 41.4
YOLOv10 [46] $\checkmark$ 7.50M 7.80ms 49.8 87.0 45.5 73.4 87.9 90.4 72.3 50.8 46.2
YOLOv11 [16] $\checkmark$ 9.74M 6.50ms 46.9 87.2 44.6 71.1 88.8 89.2 71.3 48.6 44.7
LSKNet [24] $\checkmark$ 21.8M 20.4ms 33.4 80.2 28.4 69.3 76.8 51.2 56.6 27.1 30.4
YOLOv12 [42] $\checkmark$ 9.36M 13.0ms 46.4 87.7 43.6 70.8 90.0 90.7 71.5 49.2 45.4
CFT [36] $\checkmark$ $\checkmark$ 53.8M 40.6ms 75.8 92.5 61.3 91.6 90.3 96.3 84.6 68.9 59.9
COMO [27] $\checkmark$ $\checkmark$ 22.6M 37.1ms 72.3 90.8 60.4 88.7 90.5 96.3 83.2 66.3 57.9
CLANet [14] $\checkmark$ $\checkmark$ 48.2M 29.1ms 74.8 92.2 60.7 91.3 91.6 97.2 84.6 68.5 59.6
CSSA [2] $\checkmark$ $\checkmark$ 13.5M 12.3ms 73.3 91.7 59.3 88.9 91.6 95.8 83.4 66.4 58.0
CMADet [39] $\checkmark$ $\checkmark$ 41.5M 46.7ms 70.9 90.7 52.0 86.4 91.7 97.1 81.5 63.5 55.7
ICAFusion [38] $\checkmark$ $\checkmark$ 29.0M 23.6ms 74.7 91.9 60.9 91.0 91.8 96.7 84.5 67.3 58.8
MMIDet [60] $\checkmark$ $\checkmark$ 53.8M 41.9ms 74.9 92.6 61.1 91.7 91.4 97.0 84.8 68.6 59.8
E2E-MFD [63] $\checkmark$ $\checkmark$ 31.3M 37.1ms 76.1 91.9 61.1 91.8 91.3 97.2 84.9 69.5 60.5
E2E-OSDet (Ours) $\checkmark$ $\checkmark$ 24.7M 20.9ms 77.7 90.7 64.3 91.8 92.1 97.8 85.7 70.3 61.4
IV-B4 Area-attention Fusion Module
To enhance the discriminative capacity of fused optical-SAR features across spatially diverse regions, we propose the Area-attention Fusion Module (AFM), a novel mechanism that emphasizes salient local structures and amplifies responses in critical target regions. AFM exploits the complementary characteristics of optical and SAR modalities, enabling robust and adaptive feature integration for multi-source object detection. Inspired by the efficient spatial modeling techniques in [42, 8], AFM partitions the input feature map of resolution $(H,W)$ into $k$ non-overlapping blocks of size $(H/k,W)$ or $(H,W/k)$, without requiring explicit window segmentation. This lightweight partitioning is implemented via a simple reshaping operation, preserving spatial continuity while substantially improving computational efficiency. The fused feature representation is computed as:
$F^{fus}_{j}=\mathbb{AFM}(F^{\prime}_{Oj},F^{\prime}_{Sj}),\quad j\in\{3,4,5\},$ (10)
IV-C Loss Function
In the proposed E2E-OSDet framework, we employ a composite loss function comprising Binary Cross-Entropy (BCE) and Probabilistic IoU (ProbIoU) [29] to jointly optimize classification and localization. The overall loss is defined as:
V M4-SAR Benchmark
V-A Setups and Implementation Details
[…]
We also evaluate single-source baselines implemented via the MMRotate toolkit [35], including Rotated FCOS [43], Rotated ATSS [65], Oriented R-CNN [56], Oriented RepPoints [22], RTMDet [32], PSC [58], and LSKNet [24]. These methods use a ResNet-50 backbone [13] pre-trained on ImageNet, and are optimized using AdamW [30] (learning rate: 2e-4, weight decay: 0.05) for 72 epochs with a batch size of 32 and the same $\times512$ input resolution. During evaluation, we adopted the standard COCO evaluation protocol to measure Average Precision (AP), including $AP_{50}$, $AP_{75}$, and $mAP$.
Figure 6: (a) Effect of different handcrafted features. (b) Effect of different scanning mechanisms. (c) Effect of varying ratios of handcrafted features. (d) Effect of different attention regions. (e) Performance of E2E-OSDet across various detectors. (f) Training time of the single-source method (using SAR images). (g) Training time of the multi-source method (using optical and SAR images).
TABLE III: Ablation analysis of E2E-OSDet on M4-SAR dataset.
[…]
@@ V-A Setups and Implementation Details
V-B Main Results
We first compared multi-source fusion methods (such as CFT, CLANet, CSSA, CMADet, ICAFusion, MMIDet, and E2E-MFD) with single-source detection methods (such as R-FCOS, R-ATSS, Roi-Trans, O-RepPoint, RTMDet, and the YOLO series). Experimental results (see Table II) indicate that detection accuracy using SAR images alone is significantly lower than when using optical images alone. This gap primarily stems from the unique imaging mechanism of SAR: even when targets are visible in the image, SAR struggles to accurately estimate the shape of target bounding boxes.
In contrast, all multi-source fusion methods outperform their corresponding single-modal baselines, clearly demonstrating that cross-modal fusion effectively leverages the complementarity between SAR and optical data to significantly reduce false positives and enhance detection accuracy. Among these, our proposed E2E-OSDet achieves the highest performance across all fusion methods, with a $mAP$ of 61.4%, fully validating the importance of mitigating domain differences in cross-modal detection. Furthermore, E2E-OSDet exhibits high efficiency with only 24.7M parameters and achieves an inference speed of 20.9ms. We also evaluated its generalization capability across different detection frameworks (see Fig. 6 (e)), demonstrating strong transferability and robustness.
@@ V-B Main Results
Figure 7: Visualization results of handcrafted features.
Figure 8: Qualitative comparison between the proposed E2E-OSDet and six fusion detection methods on the proposed M4-SAR dataset.
V-C Ablation Experiment
@@
As shown in Table III (a), we adopt CSSA [2] as our baseline model, which serves as the foundation for independently evaluating the contribution of each proposed module. The results (see Table III (a), (b), and (c)) indicate that all three modules lead to performance gains, among which the AFM module provides the most substantial improvement in detection accuracy. We further assess the interoperability of the modules. As shown in Table III (e), (f), and (g), the modules are compatible with one another, and their integration does not result in any performance degradation. Finally, we integrate all modules (see Table III (h)), resulting in additional gains in overall model performance.
V-C1 Impact of Filter Augment
@@
To assess the effectiveness of the FAM, we tested a range of conventional feature descriptors. As shown in Fig. 6 (a) and Fig. 7, the results indicate that incorporating handcrafted features significantly enhances detection performance. These results mapping pixels into a handcrafted feature space effectively reduces the distribution gap between optical and SAR data. Notably, the Grad feature yields superior performance due to its ability to narrow domain differences and capture multi-scale information.
V-C2 Impact of Scanning Mechanisms
We further investigate the effectiveness of the CMIM module under different scanning mechanisms [11]. As illustrated in Fig. 6 (b), CMIM consistently enhances detection accuracy regardless of the underlying scanning strategy. This observation highlights the robustness and adaptability of CMIM, suggesting that the proposed module can serve as a versatile component that generalizes well across diverse architectural settings rather than being restricted to a particular scanning design.
V-C3 Impact of Feature Proportion and Area Size
In addition, we investigate the influence of two key design choices: the proportion of handcrafted feature input and the adoption of area attention. As shown in Fig. 6 (c) and (d), varying these hyperparameters leads to only marginal performance differences, indicating that the overall framework is relatively insensitive to such configurations. This stability suggests that our method achieves consistent performance without requiring delicate hyperparameter tuning, further underscoring its practicality in real-world deployment.
V-D Qualitative Evaluation
We present visual comparisons of detection results from various multi-source methods. As shown in Fig. 8, the proposed E2E-OSDet consistently outperforms other methods in complex scenes. In particular, in regions with limited optical information, such as cloud-covered, snow-covered, low-light, or low-resolution areas, the incorporation of SAR data significantly enhances target localization accuracy. As illustrated in Fig. 9, our proposed E2E-OSDet produces highly concentrated and spatially aligned attention maps, accurately focusing on target regions while suppressing irrelevant background activations. These experimental results demonstrate the proposed dataset’s effectiveness in evaluating the robustness and adaptability of detection models across diverse scenarios.

### v3

@@ 4.1 Experiment Setup and Implementation Details
Figure 5: Comprehensive performance evaluation under different parameter settings.
4.2 Main Results
Experimental results (see Table 2) indicate that detection accuracy using SAR images alone is significantly lower than when using optical images alone. This gap primarily stems from the unique imaging mechanism of SAR: even when targets are visible in the image, SAR struggles to accurately estimate the shape of target bounding boxes. In contrast, all multi-source fusion methods outperform their corresponding single-modal baselines, clearly demonstrating that cross-modal fusion effectively leverages the complementarity between SAR and optical data to significantly reduce false positives and enhance detection accuracy. Among these, our proposed E2E-OSDet achieves the highest performance across all fusion methods, with a $mAP$ of 61.4%, fully validating the importance of mitigating domain differences in cross-modal detection. Furthermore, E2E-OSDet exhibits high efficiency with only 24.7M parameters and achieves an inference speed of 20.9ms. As shown in Fig. 5 (e), we also evaluated its generalization capability across different detection frameworks (YOLO series, including v5, v6, v8, v9, v10.), demonstrating strong transferability and robustness.
Table 3: Ablation analysis of E2E-OSDet on M4-SAR dataset.
  No. FAM CMIM AFM Train-Params. $\textbf{AP}_{50}$ $\uparrow$ $\textbf{AP}_{75}$ $\uparrow$ mAP $\uparrow$
@@ 4.2 Main Results
Figure 6: Visualization results of handcrafted features (Canny, WST, Haar, Hog, Grad).
Figure 7: Qualitative comparison of the proposed E2E-OSDet and six fusion methods.
4.3 Ablation Experiment
@@
As shown in Table 3 (a), we adopt CSSA [cao2023multimodal] as our baseline model, which serves as the foundation for independently evaluating the contribution of each proposed module. The results (see Table 3 (a), (b), and (c)) indicate that all three modules lead to performance gains, among which the AFM module provides the most substantial improvement in detection accuracy. We further assess the interoperability of the modules. As shown in Table 3 (e), (f), and (g), the modules are compatible with one another, and their integration does not result in any performance degradation. Finally, we integrate all modules (see Table 3 (h)), resulting in additional gains in overall model performance.
@@
Impact of Filter Augment: To assess the effectiveness of the FAM, we tested a range of conventional feature descriptors. As shown in Fig. 5 (a) and Fig. 6, the results indicate that incorporating handcrafted features significantly enhances detection performance. These results mapping pixels into a handcrafted feature space effectively reduces the distribution gap between optical and SAR data. Notably, the Grad feature yields superior performance due to its ability to narrow domain differences and capture multi-scale information.